I am having a hard time breaking down story problems, I can solve once I have the equation.
Here is my ?
Charlie and Doggie are squirrels that live in the trees in my mother's yard. Before snow covered the fall crop of acorns, Doogie gathered ten pounds more that two-thirds as many acorns as Charlie. Now, since Doogie played in the road and took himself out of the gene pool, Charlie has all thirty-five pounds of acorns they gathered to himself. How may pounds did each of them gather?
Thank You!
Beth
: I am having a hard time breaking down story problems, I can solve once I have the equation.
Here is my ?
Charlie and Doggie are squirrels that live in the trees in my mother's yard. Before snow covered the fall crop of acorns, Doogie gathered ten pounds more that two-thirds as many acorns as Charlie. Now, since Doogie played in the road and took himself out of the gene pool, Charlie has all thirty-five pounds of acorns they gathered to himself. How may pounds did each of them gather?
Thank You!
Beth

Answer by ankor@dixie-net.com(3845)

Charlie and Doggie are squirrels that live in the trees in my mother's yard. Before snow covered the fall crop of acorns,
:
Write an equation for the statements:
:
" Doogie gathered ten pounds more than two-thirds as many acorns as Charlie."
D = 2/3C + 10
:
" Charlie has all thirty-five pounds of acorns they gathered to himself."
C + D = 35
D = (35-C)
:
How may pounds did each of them gather?
Substitute (35-C) for D in the 1st equation, find C
;
Do you think you can handle it now?